### v2

@@ 4 Discussion
In the context of lawlike (computable) universes and formal learning systems, the epistemic consequences are more profound in the sense that these findings show that there is a ceiling above which formal knowledge actually cannot further decrease the probability of zemblanitous findings. It constitutes a double-edge-sword argument that may appear counter-intuitive at first glance. On the one hand, a formal theory that can only explain or predict (or is only corroborated by) a very small subset of events in a lawlike universe would be considered insufficiently valid or robust. On the other hand, a formal theory that explains too much can never neglect the possibility that it was deceived.
@@
Furthermore, we argue this epistemological limitation can have impacts on societal dynamics. Despite the advancements in technology of the past 100 years, two very different illusions on the state of technology are evident. The first, often dubbed “post-digital”, is the belief that the digital revolution is over as a consequence of current practical limitations (such as processing speed and storage space), concluding that human intelligence is ultimately superior to machine intelligence. However, these limitations are likely temporary and far from the epistemic limits of computing. The other illusion is a total, unwavering belief that the “promise” of machine learning will result in a technology that achieves and surpasses human intelligence.
@@
To understand the societal impact of developing technologies such as artificial intelligence and machine learning, a paradigmatic shift is necessary: one where society looks deeper into the interplay between humans and machines, and how they are similar and intertwined rather than how they are different. We posit a ‘meta-digital’ reexamination of these advancements that advocates for a more humanistic use of technology, but accepts that the digital revolution is still in its infancy and reflects the aforementioned paradigm shift. One in which artificial intelligence is not necessarily viewed as inferior nor superior to human intelligence, but rather a self-reflexive view which illuminates something that is actually shared by humans and computers: an epistemic limit of knowledge production itself. By confronting the property of digitally intermediated social dynamics being always prone to be trapped into zemblanitous findings, society will hopefully be compelled to think about, challenge, and reconsider the conceptual frameworks within which they understand and relate to technology. This, in turn, can lead to mechanisms that explain, predict, amplify, or mitigate certain collective digital-social dynamics.
Our society is at a tipping point. Governments and tech companies are collecting troves of personal data using unprecedented mass surveillance Zuboff (2018). People are either surveilled without consent or lured into contributing large amounts of data under the pretense that they are being provided a product/service, when in fact their data and the ability to manipulate their behavior is the product. In addition to the problem of the environmental impact of such technologies for the emissions of CO2 in the algorithm’s training processes and respective technological waste, this commodity is then used and sold for ulterior, sometimes nefarious motives such as suppressing populations and influencing politics. How the data are algorithmically controlled and sanctioned also amplifies political polarization, civil unrest, and systemic inequalities. Because polarizing content engages users, companies create algorithms that sow and reinforce divisive beliefs in order to maximize profits Lanier (2018). This practice notoriously amplifies and perpetuates discriminatory biases that disproportionately affect minorities Noble (2018); Cavassane et al. (2023).
Social networks and the underlying algorithms that intermediate social relationships not only can reflect political polarization and systemic inequality, but also can cause and amplify them. We argue this is a societal impact that results from the theoretical or epistemological limitations discussed in this article, therefore highlighting the importance of fostering the study of such limitations. While there is some effort to understand the problem, governments and corporations are not properly reckoning with the fundamental causes of these problems, nor are they incentivized to.
@@
Without a deeper understanding of the fundamental laws and nature of digitally intermediated societies, the very institutions created to protect people may become so irreconcilably affected that humanity as a whole is threatened. Despite society being at such a critical juncture, this research comes with a certain optimism. One that envisions a future where innovation can be compelled by mitigating systemic inequalities and biases and within which all members of society have more agency in the development and maintenance of a more humane and equitable political and economic framework.
@@ 5 Acknowledgments
Felipe S. Abrahão acknowledges the partial support from São Paulo Research Foundation (FAPESP), grant 2023202320232023/05593055930559305593-$11$. Ricardo P. Cavassane acknowledges the partial support from the Coordination for the Improvement of Higher Education Personnel (CAPES), Brazil, Finance Code $001001$. Mariana Vitti Rodrigues acknowledges the support from São Paulo Research Foundation (FAPESP), grant 2023202320232023/01408014080140801408-$55$.

### v3

@@ 4 Discussion
In the context of lawlike (computable) universes and formal learning systems, the epistemic consequences are more profound in the sense that these findings show that there is a ceiling above which formal knowledge (either produced by computation-assisted science or by artificial machines) actually cannot further decrease the probability of zemblanitous findings. It constitutes a double-edge-sword argument that may appear counter-intuitive at first glance. A formal theory that can only explain or predict (or is only corroborated by) a very small subset of events in a lawlike universe would be considered insufficiently valid or robust; while a formal theory that explains too much can never neglect the possibility that it was deceived.
@@
Furthermore, we argue this epistemological limitation can have impacts on societal dynamics. Despite the advancements in technology of the past 100 years, two very different misconceptions on the state of technology are evident. The first, often dubbed “post-digital”, is the belief that the digital revolution is over as a consequence of current practical limitations (such as processing speed and storage space), concluding that human intelligence is ultimately superior to machine intelligence. However, these limitations are likely temporary and far from the epistemic limits of computing. The other misconception is a total, unwavering belief that the “promise” of machine learning will result in a technology that achieves and surpasses human intelligence.
@@
Contrary to leaving the discussion and scientific research on the societal impacts of AI as a secondary problem in the everyday practice of AI developers Birhane et al. (2022), we argue that a paradigmatic shift is necessary to understand these impacts: one where society looks deeper into the interplay between humans and machines, and how they are similar and intertwined rather than how they are different. We posit a ‘meta-digital’ reexamination of these advancements that advocates for a more humanistic use of technology, but accepts that the digital revolution is still in its infancy and reflects the aforementioned paradigm shift. One in which AI is not necessarily viewed as inferior nor superior to human intelligence, but rather a self-reflexive view which illuminates such a limitation shared by humans and computers: an epistemic limit of formal knowledge production itself to which any formal learning system, be it human or artificial, are prone. This shared limitation in turn highlights the generality of the results presented in this article. By confronting the property of digitally intermediated social dynamics being always prone to be trapped into zemblanitous findings, society will hopefully be compelled to think about, challenge, and reconsider the conceptual frameworks within which they understand and relate to technology. This, in turn, can lead to mechanisms that explain, predict, amplify, or mitigate certain collective digital-social dynamics.
We believe that our society is at a crucial point in which governments and tech companies are collecting troves of personal data using unprecedented mass surveillance Zuboff (2018). People can be surveilled without consent, or willingly contribute large amounts of data by subscribing to agreement terms that are required so they can use the respective tool, application, or platform. In addition to the problem of the environmental impact of such technologies for the emissions of CO2 in the algorithm’s training processes and respective technological waste, personal data is a commodity that can be used and sold. There is a concern regarding whether or not social networks have given rise to echo chambers and filter bubbles Cinelli et al. (2021); Terren et al. (2021) that can polarize political opinions or influence digital behavior Chitra and Musco (2020); Kashima et al. (2021). Because polarizing content engages users, algorithms could be created to sow and reinforce divisive beliefs, e.g. in order to maximize profits Lanier (2018). Social networks and the underlying algorithms that intermediate social relationships can reflect political polarization and systemic inequality Lanier (2018); Noble (2018), and may also amplify them Cavassane et al. (2023); O’Neil (2016). We argue these are societal impacts that highlight the importance of fostering the study of the theoretical or epistemological limitations discussed in this article.
@@
Without a deeper understanding of the fundamental laws and nature of digitally intermediated societies, the very institutions created to protect people may become so irreconcilably affected that humanity as a whole is threatened. With the purpose of mitigating systemic inequalities and biases, we propose the fostering of scientific research in this field and the development of novel technological tools. Those new tools should enable or reinforce that all members of society have more agency in the development and maintenance of a more humane and equitable political and economic framework.
@@ 5 Acknowledgments
Felipe S. Abrahão acknowledges support from São Paulo Research Foundation (FAPESP), grants 2021202120212021/14501145011450114501-$88$ and 2023202320232023/05593055930559305593-$11$. Ricardo P. Cavassane acknowledges support from the National Council for Scientific and Technological Development (CNPq), Brazil, grant 150915150915150915150915/2024202420242024-$11$, and from the Coordination for the Improvement of Higher Education Personnel (CAPES), Brazil, Finance Code $001001$. Mariana Vitti-Rodrigues acknowledges the support from São Paulo Research Foundation (FAPESP), grant 2023202320232023/01408014080140801408-$55$.